The half-life of strontium-91, 91


38Sr is 9.70 h. Find


(a) its decay constant, and (5pts)

(b) for an initial 1.00-g sample, the activity after 15 hours.


1
Expert's answer
2021-10-29T02:30:13-0400

Half life of "^{91}_{38} Sr= T_{\frac{1}{2}}= 9.7"

= 9.7×60×60

=34920s

"T_{\frac{1}{2}}\lambda" = 0.693

"\lambda" = 0.693/T"_\frac{1}{2}"

= 0.693/34920

decay constant= 1.9845×10-5